# Network Modelling of Asynchronous Change-Points in Multivariate Time Series

### v1

@@ Acknowledgements
Sampling the hidden states in series $j$, $\boldsymbol{x}_{j,0:T}$ using algorithm 1 requires additional algorithms to perform resampling and backward sampling. Since we are performing conditional particle filtering within a particle Gibbs framework, the previously sampled hidden states for series $j$ are guaranteed to survive all resampling steps when filtering. We therefore use the conditional SOR algorithm described by algorithm 2. All of the below algorithms can be found in (Whiteley et al.,, 2011; Fearnhead and Liu,, 2007).
@@
1:$(\mathcal{S}_{t},\mathcal{W}_{t})$ particles/weights to resample, $x_{j,t}^{*}$ (current hidden state), $N$ (number particles to keep).
@@
3:For each $s\in\mathcal{S}_{t}$, if $w_{t}(s)>\kappa$, place particle $s$ into set $A$. Otherwise place $s$ into set $B$.
4:Let $(\mathcal{S}_{t}^{A},\mathcal{W}_{t}^{A})$ and $(\mathcal{S}_{t}^{B},\mathcal{W}_{t}^{B})$ be the particles & weights for sets $A$ and $B$ respectively.
@@
7:     Sample $N-K$ particles from $(\mathcal{S}_{t}^{B},\mathcal{W}_{t}^{B})$ using the Conditional Stratified Resampling Algorithm (conditional on $x_{j,t}^{*}$).
@@
9:     Sample $N-K$ particles from $(\mathcal{S}_{t}^{B},\mathcal{W}_{t}^{B})$ using the Stratified Resampling Algorithm.
@@
1:$(\mathcal{S}_{t}^{B},\mathcal{W}_{t}^{B})$ particles/weights to resample, $x_{j,t}^{*}$ (current hidden state), $N-K$ (number particles to keep).
2:Normalise the weights, $\mathcal{W}_{t}^{B}$, such that $\sum_{s\in\mathcal{S}_{t}^{B}}w_{t}^{B}(s)=1$.
@@
4:Sample $V^{*}$ uniformly on $[Q_{t}(x_{j,t}^{*}-1),Q_{t}(x_{j,t}^{*})]$, set $V_{1}=V^{*}-\frac{\lfloor(N-K)V^{*}\rfloor}{N-K}$ and $V_{p}=V_{p-1}+1/(N-K)$ for $p=2,...,N-K$.
@@
1:$(\mathcal{S}^{B}_{t},\mathcal{W}^{B}_{t})$ particles/weights to resample, $N-K$ (number particles to resample).
2:Normalise the weights, $\mathcal{W}_{t}^{B}$, such that $\sum_{s\in\mathcal{S}_{t}^{B}}w_{t}^{B}(s)=1$.
@@
After generating the particle approximations for each time $t=1,...,T$, we generate a new sample $\boldsymbol{x}_{j,0:T}$ using backward sampling given in Algorithm 5.
@@
1:$\{(\mathcal{S}_{t},\mathcal{W}_{t})\}_{t=1,...,T}$ particle/weight sets after performing particle filtering on series $j$, $\boldsymbol{x}^{*}$ (current matrix of hidden states).
@@
We sample $\boldsymbol{A}$ by Gibbs sampling each pair $(A_{i,j},A_{j,i})$ for $i=1,...,d,j=i+1,...,d$. The full posterior conditional for each pair $(A_{i,j},A_{j,i})$ is proportional to:
@@
×$\times\prod_{t=1}^{t}Pr(X_{i,t}=x_{i,t}|\boldsymbol{X}_{t-1}$ $=\boldsymbol{x}_{t-1},\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A}^{*},\boldsymbol{q}_{0},\boldsymbol{q})Pr(X_{j,t}=x_{j,t}|\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1},\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A}^{*},\boldsymbol{q}_{0},\boldsymbol{q})$ (38)
defined for $(a_{i,j},a_{j,i})\in\{(1,0),(0,1),(0,0)\}$ and $\boldsymbol{A}^{*}$ is the current adjacency matrix with the $a_{i,j}$ and $a_{j,i}$ imputed at $(i,j)$ and $(j,i)$ respectively. Thus, $(A_{i,j},A_{j,i})$ can be updated using a categorial sample.
@@
Due to correlation between $(W_{i,j},q_{i,j})$, we find that it is beneficial to update these jointly for $j=1,...,d;\;i=0,...,d$. We update $(W_{i,j},q_{i,j})$ with a random walk Metropolis step using independent Gaussian proposals: $W_{i,j}^{\dagger}\sim\mathcal{N}(W_{i,j},0.5^{2})$ and $q_{i,j}^{\dagger}\sim\mathcal{N}(q_{i,j},0.05^{2})$. We then accept $(W_{i,j}^{\dagger},q_{i,j}^{\dagger})$ with probability
@@
Appendix B: Comparison of Gibbs Sampling Schemes
@@
The estimated marginal posterior change-point probabilities for series one under each sampling method are shown in Figure 7. We can see that the estimated probabilities for each method are virtually identical. In order to assess the mixing properties of the methods, we may inspect the sampled hidden states at each iteration. From Figure 7 we can see that in series one, time point 583 has around 0.6 probability of being included as a change-point. We can assess the mixing of $X_{1,583}$ by inspecting the sampled values of $U_{1,583}=I(X_{1,583}=582)$. I.e. $U_{1,583}=1$ for a change-point and $U_{1,583}=0$ for no change-point.
@@
From Figure 8 we can see that under the single-site Gibbs sampler, $U_{1,583}$ mixes very poorly, getting stuck in a given state for long periods of time. However, the particle Gibbs methods with 50 and 100 particles both appear to be mixing much better. This is evident in the autocorrelation plots in Figure 8. Here the single-site Gibbs exhibits a slow decay in the correlation between samples while the particle Gibbs samples decorrelate almost immediately. We then estimated the integrated autocorrelation time (IAT) of each method using the R function (“IAT”). The single-site Gibbs obtains an IAT of 58.5 while the particle Gibbs schemes both obtain IATs of 1.2. This suggests that we must obtain around 49 samples from the single-site sampler in order to obtain the same amount of information as one sample from particle Gibbs samplers (when estimating $\mathbb{E}[U_{1,583}|\boldsymbol{y}]$). Similar results can be obtained by inspecting the change-point indicators at other time points. These results suggest that relatively few particles are required to achieve good mixing, with the particle Gibbs sampler obtaining the same IAT for 50 and 100 particles.
@@
Appendix C: Marginal Segment Likelihood for Gaussian Mean Model

### v2

@@ Acknowledgments
Sampling the hidden states in series $j$, $\boldsymbol{x}_{j,0:T}$ using Algorithm 1 requires additional algorithms to perform resampling and backward sampling. Since we are performing conditional particle filtering within a particle Gibbs framework, the previously sampled hidden states for series $j$ are guaranteed to survive all resampling steps when filtering. We therefore use the conditional SOR Algorithm described by Algorithm 2. All of the below algorithms can be found in [90, 64].
@@
[…]
@@
3:For each $s\in\mathcal{S}_{t}$, if $w_{t}(s)>\kappa$, place particle $s$ into set AA. Otherwise place $s$ into set $B$.
4:Let $(\mathcal{S}_{t}^{A},\mathcal{W}_{t}^{A})$ and $(\mathcal{S}_{t}^{B},\mathcal{W}_{t}^{B})$ be the particles & weights for sets AA and $B$ respectively.
@@
7:  Sample NKN-K particles from $(\mathcal{S}_{t}^{B},\mathcal{W}_{t}^{B})$ using the Conditional Stratified Resampling Algorithm (conditional on $x_{j,t}^{*}$).
@@
9:  Sample NKN-K particles from $(\mathcal{S}_{t}^{B},\mathcal{W}_{t}^{B})$ using the Stratified Resampling Algorithm.
@@
[…]
2:Normalize the weights, $\mathcal{W}_{t}^{B}$, such that $\sum_{s\in\mathcal{S}_{t}^{B}}w_{t}^{B}(s)=1$.
@@
4:Sample $V^{*}$ uniformly on $Q_{t}(x_{j,t}^{*}-1),Q_{t}(x_{j,t}^{*})]$, set $V_{1}=V^{*}-\frac{\lfloor(N-K)V^{*}\rfloor}{N-K}$ and $V_{p}=V_{p-1}+1/(N-K)$ for p=2,,NKp=2,...,N-K.
@@
[…]
2:Normalize the weights, $\mathcal{W}_{t}^{B}$, such that $\sum_{s\in\mathcal{S}_{t}^{B}}w_{t}^{B}(s)=1$.
@@
After generating the particle approximations for each time t=1,,Tt=1,...,T, we generate a new sample $\boldsymbol{x}_{j,0:T}$ using backward sampling given in Algorithm 5.
@@
[…]
@@
We sample $\boldsymbol{A}$ by Gibbs sampling each pair $(A_{i,j},A_{j,i})$ for i=1,,d,j=i+1,,di=1,...,d,j=i+1,...,d. The full posterior conditional for each pair $(A_{i,j},A_{j,i})$ is proportional to:
@@
$\times\prod_{t=1}^{t}Pr(X_{i,t}=x_{i,t}|\boldsymbol{X}_{t-1}$ $=\boldsymbol{x}_{t-1},\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A}^{*},\boldsymbol{q}_{0},\boldsymbol{q})Pr(X_{j,t}=x_{j,t}|\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1},\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A}^{*},\boldsymbol{q}_{0},\boldsymbol{q})$ (42)
defined for $(a_{i,j},a_{j,i})\in\{(1,0),(0,1),(0,0)\}$ and $\boldsymbol{A}^{*}$ is the current adjacency matrix with the $a_{i,j}$ and $a_{j,i}$ imputed at $(i,j)$ and $(j,i)$ respectively. Thus, $(A_{i,j},A_{j,i})$ can be updated using a categorical sample.
@@
Due to correlation between $(W_{i,j},q_{i,j})$, we find that it is beneficial to update these jointly for j=1,,d;i=0,,dj=1,...,d;\;i=0,...,d. We update $(W_{i,j},q_{i,j})$ with a random walk Metropolis step using independent Gaussian proposals: $W_{i,j}^{\dagger}\sim\mathcal{N}(W_{i,j},0.5^{2})$ and $q_{i,j}^{\dagger}\sim\mathcal{N}(q_{i,j},0.05^{2})$. We then accept $(W_{i,j}^{\dagger},q_{i,j}^{\dagger})$ with probability
@@
Appendix C: Comparison of Gibbs Sampling Schemes
@@
The estimated marginal posterior change-point probabilities for series one under each sampling method are shown in Figure 9. We can see that the estimated probabilities for each method are virtually identical. In order to assess the mixing properties of the methods, we may inspect the sampled hidden states at each iteration. From Figure 9 we can see that in series one, time point 583 has around 0.6 probability of being included as a change-point. We can assess the mixing of $X_{1,583}$ by inspecting the sampled values of $U_{1,583}=\boldsymbol{1}(X_{1,583}=582)$. I.e. $U_{1,583}=1$ for a change-point and $U_{1,583}=0$ for no change-point.
@@
From Figure 10 we can see that under the single-site Gibbs sampler, $U_{1,583}$ mixes very poorly, getting stuck in a given state for long periods of time. However, the particle Gibbs methods with 50 and 100 particles both appear to be mixing much better. This is evident in the autocorrelation plots in Figure 10. Here the single-site Gibbs exhibits a slow decay in the correlation between samples while the particle Gibbs samples decorrelate almost immediately. We then estimated the integrated autocorrelation time (IAT) of each method using the R function “IAT”. The single-site Gibbs obtains an IAT of 58.5 while the particle Gibbs schemes both obtain IATs of 1.2. This suggests that we must obtain around 49 samples from the single-site sampler in order to obtain the same amount of information as one sample from particle Gibbs samplers (when estimating $\mathbb{E}[U_{1,583}|\boldsymbol{y}]$). Similar results can be obtained by inspecting the change-point indicators at other time points. These results suggest that relatively few particles are required to achieve good mixing, with the particle Gibbs sampler obtaining the same IAT for 50 and 100 particles.
@@
Appendix D: Marginal Segment Likelihood for Gaussian Mean Model